# Inference-Aware Prompt Optimization for Aligning Black-Box Large Language Models

### v1

@@ Abstract
Prompt optimization methods have demonstrated significant effectiveness in aligning black-box large language models (LLMs). In parallel, inference scaling strategies such as Best-of-N Sampling and Majority Voting have also proven to enhance alignment and performance by trading off computation. However, existing prompt optimization approaches are inference strategy agnostic; that is, they optimize prompts without regard to the inference strategy employed during deployment. This constitutes a significant methodological gap, as our empirical and theoretical analysis reveals a strong interdependence between these two paradigms. Moreover, we find that user preferences regarding trade-offs among multiple objectives and inference budgets substantially influence the choice of prompt and inference configuration. To address this gap, we introduce a unified novel framework named Iapo (Inference-Aware Prompt Optimization) that jointly optimizes the prompt and inference scale, while being aware of the inference budget and different task objectives. We then develop a fixed-budget training algorithm for Iapo, which we call Psst (Prompt Scaling via Sequential Trimming), and analyze finite-budget guarantees on error probability. Finally, we evaluate the effectiveness of Psst on six different tasks, including multi-objective text generation and reasoning, and demonstrate the critical role of incorporating inference-awareness when aligning black-box LLMs through prompt optimization.
@@ Introduction
In recent years, most state-of-the-art large language models (LLMs) are accessible only through black-box APIs. Traditional alignment methods that require access to model weights or logits are therefore infeasible. To address this issue, prompt optimization-based alignment methods have garnered interest (Chang et al. 2024). These methods typically enhance input prompts by rewording or appending additional instructions to better align the models’ outputs with a task’s objectives. Another broadly applicable alignment strategy for black-box models is scaling inference computations using strategies such as Best-of-N Sampling or Majority Voting. These inference scaling methods generate multiple candidate responses for the same query and select the final response via ranking or voting mechanisms (Wang et al. 2022; Krishna et al. 2022; Gui, Gârbacea, and Veitch 2024; Yue et al. 2025).
@@
Although existing prompt optimization techniques have achieved substantial success, they are typically agnostic to how model outputs are aggregated or sampled, overlooking the impact of such inference methods. Our initial empirical investigation reveals that the performance of optimized prompts is highly sensitive to the choice of inference-scaling approach. Furthermore, our theoretical analysis reveals that decoupling prompt optimization from inference can lead to misalignment. Finally, we observe that optimal alignment requires careful consideration of user-specific preferences regarding the trade-offs among multiple objectives, as well as the computational resources they are willing to expend. These findings expose a critical gap in current methods: the absence of a unified framework that simultaneously accounts for prompt optimization, inference-scaling strategies, user preferences, and computational resource constraints.
@@
To bridge this gap, we introduce Iapo (Inference-Aware Prompt Optimization), a novel prompt optimization framework designed explicitly to produce aligned responses from inference-scaled black-box LLMs. Iapo simultaneously optimizes prompt design and inference scaling strategies while considering different task objectives and computational budgets. We formulate the task of identifying an optimal policy for the Iapo framework as a contextual best-arm identification (BAI) problem. To efficiently solve this, we propose a fixed-budget training algorithm named Psst (Prompt Scaling via Sequential Trimming). Additionally, we introduce a warm-up heuristic that further improves performance within the training budget.
@@
We begin our analysis by deriving theoretical finite-budget guarantees on the error probability of Psst. Next, we empirically demonstrate the effectiveness of Psst for learning Iapo policies across six diverse tasks, including multi-objective text generation, mathematical reasoning, and commonsense reasoning benchmarks. Additionally, our analysis shows that ignoring inference scaling during prompt optimization can lead to substantial misalignment, highlighting the critical role of inference-awareness in aligning black-box LLMs.
@@
Over the years, considerable effort has been devoted to aligning large language models (LLMs) with human expectations in downstream tasks (Minaee et al. 2024). Many widely adopted alignment approaches—such as Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert 2025)—require access to model weights. This limitation has motivated increasing interest in black-box alignment methods such as prompt optimization, which can align black-box models through input manipulation alone (Zhou et al. 2022; Ouyang et al. 2022; Chang et al. 2024). Prompt optimization has demonstrated strong performance in both single-objective (Cheng et al. 2023; Trivedi et al. 2025) and multi-objective (Jafari et al. 2024; Zhao et al. 2025) settings. However, these methods are agnostic of the inference strategy used during deployment, potentially leading to suboptimal performance. In contrast, our approach explicitly addresses the interdependence of inference-time strategy and prompt optimization.
@@
Recently, Shi et al. framed prompt optimization as a fixed-budget best-arm identification (BAI) problem. While effective under limited evaluation budgets, the method remains inference agnostic and was only explored in single-objective settings. Our work builds on this foundation in two key ways: (1) we introduce a contextual formulation that models user preferences over multiple objectives and associated computational costs; and (2) we incorporate inference-awareness to ensure alignment with the actual inference strategy. To learn an optimal policy, we introduce a fixed-budget contextual BAI algorithm, Psst, inspired by Sequential Halving (SH) (Karnin, Koren, and Somekh 2013). While SH was originally developed for the pure bandit setting, the Iapo framework features both inter-context full-information feedback and intra-context semi-bandit feedback. Psst leverages these structural properties to achieve more efficient optimization, extending beyond what standard SH can accommodate.
@@
Another relevant line of work focuses on inference-time alignment, where model outputs are improved during inference without modifying model parameters. Some of these methods, such as GenARM and DEAL (Xu et al. 2024; Huang et al. 2024), require access to model logits, limiting their applicability in black-box settings. In contrast, Best-of-N sampling (BoN) and Majority Voting (MV) methods operate purely on model outputs and have shown strong empirical gains by generating multiple candidates and selecting the best one (OpenAI 2024; Yue et al. 2025; Wang et al. 2022; Krishna et al. 2022). However, these approaches introduce a non-trivial computational cost, and to our knowledge, none of them explicitly optimize the trade-off between computational budget and output quality. Further, our preliminary experiments show that such inference-scaling strategies interact non-trivially with prompt design: prompts optimized for single-shot decoding may yield suboptimal performance under BoN or MV, and vice versa. This necessitates an inference-aware prompt optimization framework.
@@
Finally, some white-box methods have recently integrated inference-awareness into the training process. For example, Chow et al. (2025) proposed an inference-aware fine-tuning procedure that explicitly optimizes for exploration–exploitation trade-offs under BoN. Similarly, BOND (Sessa et al. 2024) and BonBon (Gui, Gârbacea, and Veitch 2024) aim to distill BoN policies into a single-pass generation process through supervised fine-tuning. While these approaches avoid the cost of sampling at inference time, they require full access to model parameters and do not generalize beyond BoN-style strategies. In contrast, our method is complementary to inference-aware fine-tuning designed for black-box LLM.
Inference–Aware Prompt Optimization
@@
Let $\mathcal{X}$ be the distribution of user queries and $\mathcal{P}$ a finite prompt set. A pair $(x\in\mathcal{X},p\in\mathcal{P})$ is submitted to a frozen black-box LLM, which, under fixed decoding hyperparameters, generates $N\in\{1,\dots,N_{\max}\}$ i.i.d. completions $\mathbf{y}_{1:N}=(y_{1},\dots,y_{N})$. $K$ bounded (potentially vector) objectives score each completion $O_{k}:\mathcal{X}\times\mathcal{Y}\to[o_{k}^{\min},o_{k}^{\max}]$ (e.g. helpfulness, harmlessness, exact-match). We also define the cost of producing a response as $\mathrm{Cost}(x,y_{i})$, a composite function that takes into account various computational factors such as token count, time, and energy. We add it as a (K+1)(K{+}1)-st objective $O_{k+1}=-\mathrm{Cost}(x,y_{i})$. An external entity supplies a context $c=(w_{1},\dots,w_{K+1})\in\mathcal{C}$, where every $w_{k}$ is chosen from a finite discrete domain. Given the above setup, we formalize the inference strategy as follows.
@@
$R_{x}^{\textsc{BoN}}(c,p,N)=\underbrace{\max_{i\leq N}\sum_{k=1}^{K}w_{k}\,o_{k}\bigl{(}x,y_{i}\bigr{)}}_{\text{task reward}}+\underbrace{w_{k+1}\sum_{i=1}^{N}o_{k+1}(x,y_{i})}_{\text{inference cost}}.$ (1)
@@
For query $x$, the pair $(p,N)$ yields i.i.d. completions $y_{1:N}$ and extracted answers $\ell_{i}=h(x,y_{i})$. For each distinct answer $s$, define the vote count $n_{s}=\sum_{i=1}^{N}\mathbf{1}[\ell_{i}=s]$, the maximum $n^{\star}=\max_{s}n_{s}$, and the tie multiplicity $t=\sum_{s}\mathbf{1}[n_{s}=n^{\star}]$. MV predicts uniformly at random among the $t$ maximizers. With gold answer $a(x)$ and the success credit defined as $o_{1}(x,p,N)=\frac{\mathbf{1}[\,n_{a(x)}=n^{\star}\,]}{t}$ we define MV utility as:
$R_{x}^{\textsc{MV}}(c,p,N)=\underbrace{w_{1}\,o_{1}(x,p,N)}_{\text{task reward}}\;+\;\underbrace{w_{2}\sum_{i=1}^{N}o_{2}(x,y_{i},c)}_{\text{inference cost}}.$ (2)
@@
(a) MATH, Majority Vote
(b) Helpful-Harmless, Best-of-N (Prompt-A)
(c) Helpful-Harmless, Best-of-N (Prompt-B)
Figure 1: Prompt–Inference Interdependence. (a) Accuracy under Majority Voting with LLaMA-3.3-70B-Instruct, showing prompt dominance shifts with budget (shaded). (b, c) Cost-adjusted reward under Best-of-N decoding. Prompt and inference scales vary with user-defined trade-offs.
@@
Let an inference configuration be a tuple $\theta\in\Theta$ (e.g. temperature, top-$p$, max token). Then we define a set of arms $\mathcal{A}$ in Iapo as: $a\;=\;(p,\theta,N)\;\in\;\mathcal{A}:=\mathcal{P}\times\Theta\times\{1,\dots,N_{\max}\}.$
@@
Given a dataset $\mathcal{X}$, context $c\in\mathcal{C}$, and aggregator $\alpha\in\{\textsc{BoN},\textsc{MV}\}$ the expected utility of arm $a$, i.e., the context-action value function or $Q$-function is defined as:
@@
Note that $R_{x}^{\alpha}(c,a)$ is a random variable. Now, let the context-optimal arm is $a^{\star}(c)=\arg\max_{a}Q^{\alpha}(a,c)$; hence the optimal Iapo policy is defined as: $\pi^{\star}(c)\;=\;a^{\star}(c),\forall\,c\in\mathcal{C}$.
In this paper, we adopt a train-then-deploy setup to learn the optimal IAPO policy. Given a total completion budget of $T$, the learner may adaptively select arms $a_{t}=(p_{t},N_{t})\in\mathcal{A}$ and query $x_{t}\sim\mathcal{X}$, then observe the full raw reward vector $\mathbf{m}_{t}\in\mathbb{R}^{K+1}$ for all completions. This process may continue until the budget is exhausted ($\sum_{t}N_{t}=T$). After spending the entire budget, the learner returns a deployment policy $\pi_{T}$. The performance of this policy is evaluated by the Average Contextual Return:
$\text{ACR}(\pi_{T})\;=\;\mathbb{E}_{c\sim\mathcal{C}}\bigl{[}Q^{\alpha}(c,a)\bigr{]},$ (4)
@@
To see how the green and blue trend can emerge, consider the following example. Suppose in a reasoning task with MV, for Prompt 1 we have $40\%$ in Query 1, $90\%$ in Query 2, and for Prompt 2, $62\%$ (both queries). Single-shot averages favor A ($0.65$ vs. $0.62$), but under MV with $N=10$, A drops to $\approx 0.63$ while B improves to $\approx 0.77$.
(b,c) Best-of-$\boldsymbol{N}$ on Helpful–Harmless. We evaluate two prompts (A and B) on the Helpful–Harmless benchmark (Bai et al. 2022) using Best-of-$N$ decoding for $N\leq 24$. Each curve corresponds to a different user-defined trade-off between helpfulness and harmlessness, plotting the cost-adjusted reward averaged over 1000 queries (see the Appendix for details). The optimal choice of prompt (A vs. B) and sampling budget (N) is highly sensitive to these preferences. For example, the prompt A is strictly preferred when helpfulness is weighted more heavily.
Having established the need for inference-aware optimization, we now examine the precise conditions under which joint optimization becomes essential. We start by establishing the Inference-Agnostic (IA) utility:
@@
$R_{x}^{\textsc{IA}}(c,a=(p,N))=\frac{1}{N}\textstyle\sum_{i=1}^{N}\sum_{k=1}^{K}w_{k}o_{k}(x,y_{i}).$ (5)
@@
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $kR_{x}^{\textsc{IA}}(c,a),k>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$Q^{AF}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl{[}aR_{x}^{IA}(c,a)+b\bigr{]}=kQ^{IA}(c,a)+b.\quad$
Figure 2: Expected utility ($w_{k+1}=0$) for MV (left) and BoN (right). MV shows a sharp performance drop when the correctness probability drops below 0.5, whereas BoN is strictly concave.
The above also highlights that affine aggregation significantly simplifies inference-aware optimization. For instance, in a regression task where the aggregated prediction is the mean of multiple numeric predictions and the reward is defined by the mean squared error (MSE), in some cases can become an affine transformation of IA, eliminating the need to simulate inference scaling during training. However, common inference scaling strategies like BoN and MV generally do not admit such affine formulations. While they can sometimes be expressed as non-affine transformations of the IA—such as in the Bernoulli case with large $N$, where $R_{x}^{\textsc{IA}}(c,a)\approx p$ (Figure 2)—these are special cases. Hence, trying to determine the prompt based on $Q^{IA}$ for BoN or MV will result in misalignment. This motivates the next section, where we develop a training method that handles the general Iapo setting beyond the affine regime.
@@
In this section, we propose a fixed-budget arm elimination-based strategy for training policy $\pi_{T}$, called Psst (Prompt Scaling via Sequential Trimming). We then provide a theoretical analysis that establishes error guarantees for Psst under a finite inference budget. Finally, we introduce a practical approximation heuristic that improves computational efficiency without significantly compromising performance in many practical settings.
@@
Classical arm–elimination methods such as Sequential Elimination (Even-Dar, Mannor, and Mansour 2006) and Sequential Halving (Karnin, Koren, and Somekh 2013) follow a simple recipe: (i) split the elimination process into multiple rounds; (ii) in each round, allocate the round budget across the surviving arms; and (iii) trim a subset of arms at the end of the round based on their estimates. However, Iapo departs from pure BAI settings in the following key ways:
@@
  • Cross‑context reuse. One pull of $(p,N)$ on query $x$ yields the completion set $\mathbf{y}_{1:N}$ and objective vector set $\mathbf{o}_{1:N}$ that can can be used to estimate $R^{\alpha}_{x}(c,p,N)$ for all $c\in\mathcal{C}$.
  • Nested sample reuse across inference scales. Pulling a larger scale subsumes smaller ones: a pull of $(p,N_{i})$ produces $\big{\lfloor}N_{i}/N_{j}\big{\rfloor}$ i.i.d. block samples for arm $(p,N_{j})$ by partitioning the $N_{i}$ draws into disjoint groups of size $N_{j}$ (e.g., to recompute BoN/MV on each block).
A key consequence is that, for a prompt, the largest surviving scale drives the budget. Let $N^{(r)}_{\max}(p)=\max\{\,N:\ (p,N)\ \text{survives at the start of round }r\,\}.$ If we allocate $K$ pulls (blocks) to $(p,N^{(r)}_{\max}(p))$ in round $r$, then every surviving arm $(p,N)$ with $N\leq N^{(r)}_{\max}(p)$ automatically receives at least $K$ effective samples by block reuse. Thus, an effective arm elimination strategy should exploit both (i) cross‑scale reuse within prompts and (ii) cross‑context reuse when scoring, while being aware of asymmetric cost.
@@
1:Context set $\mathcal{C}$, prompt set $\mathcal{P}$, scale set $\mathcal{N}$, Scaling strategy $\alpha$, Query Dataset $\mathcal{X}$, total pull budget $T$;
@@
7:  $n_{r}\leftarrow\bigl{\lfloor}T/R\bigr{\rfloor}$
8:  $\lambda^{(r)}\leftarrow\textsc{Allocate}\bigl{(}\text{F},n_{r}\bigr{)}$
9:  $\mathcal{B}\leftarrow\{\}$
10:  for $(a,n_{r})\in\lambda^{(r)}$ do
11:   for $i=1...n_{r}$ do
12:     Sample $x\sim\mathcal{X}$
13:     $\mathcal{B}\leftarrow\mathcal{B}\ \cup(a,x)$
14:   end for

### v2

@@ Abstract
Prompt optimization methods have demonstrated significant effectiveness in aligning black-box large language models (LLMs). In parallel, inference scaling strategies such as Best-of-N Sampling and Majority Voting have likewise been shown to improve alignment and performance by trading additional computation for better output. However, existing prompt optimization approaches are inference strategy agnostic; that is, they optimize prompts without accounting for the inference strategy. This constitutes a significant methodological gap, as our empirical and theoretical analysis reveals a strong interdependence between these two paradigms. Moreover, we find that user preferences regarding trade-offs among multiple objectives and inference budgets substantially influence the choice of prompt and inference configuration. To address this gap, we introduce a novel unified framework named Iapo (Inference-Aware Prompt Optimization) that jointly optimizes the prompt and inference scale, while being aware of the inference budget and different task objectives. We then develop a fixed-budget training algorithm for Iapo, called Psst (Prompt Scaling via Sequential Trimming), and establish finite-budget guarantees on the error probability. Finally, we evaluate the effectiveness of Psst on six tasks, including multi-objective text generation and reasoning, and demonstrate the critical role of incorporating inference-awareness in aligning black-box LLMs using prompt optimization.
@@ Introduction
Most state-of-the-art large language models (LLMs) are currently accessible exclusively through black-box APIs. Traditional alignment methods that require access to model weights or logits are therefore infeasible. To address this challenge, prompt-based alignment methods have gained substantial attention in recent work (Chang et al. 2024). These methods typically enhance input prompts by rewording them or appending additional instructions to better align the models’ outputs with a task’s objectives. Another broadly applicable alignment method for black-box models is scaling inference computations using strategies such as Best-of-N sampling or Majority Voting. These inference scaling methods generate multiple candidate responses for the same query and select the final response via ranking or voting mechanisms (Krishna et al. 2022; Wang et al. 2022; Gui et al. 2024; Yue et al. 2025).
@@
Although existing prompt optimization techniques have achieved substantial success, they are typically agnostic to how model outputs are aggregated or sampled, overlooking the impact of such inference methods. Our initial empirical investigation reveals that the performance of optimized prompts is highly sensitive to the choice of inference scaling approach. Furthermore, our theoretical analysis reveals that decoupling prompt optimization from inference can lead to misalignment. Finally, we observe that optimal alignment requires careful consideration of user-specific preferences regarding the trade-offs among multiple objectives, as well as the computational resources users are willing to expend. These findings expose a critical gap in current methods: the absence of a unified framework that simultaneously accounts for prompt optimization, inference scaling strategies, user preferences, and computational resource constraints.
@@
To bridge this gap, we introduce Iapo (Inference-Aware Prompt Optimization), a novel prompt optimization framework designed explicitly to produce aligned responses from inference-scaled black-box LLMs. Iapo simultaneously optimizes prompt design and inference scaling strategies while considering different task objectives and computational budgets (Figure 1). We formulate the task of identifying an optimal policy for the Iapo framework as a contextual best-arm identification problem. To efficiently solve this, we propose a fixed-budget training algorithm named Psst (Prompt Scaling via Sequential Trimming). Additionally, we introduce a warm-up heuristic that further improves performance within the training budget.
@@
We begin our analysis by deriving theoretical finite-budget guarantees on the error probability of Psst. Next, we empirically demonstrate the effectiveness of Psst for learning Iapo policies across six diverse tasks, including multi-objective text generation, mathematical reasoning, and commonsense reasoning benchmarks. Additionally, our analysis shows that ignoring inference scaling during prompt optimization can lead to substantial misalignment, highlighting the critical role of inference-awareness in aligning black-box LLMs. The results establish that prompt quality cannot be decoupled from the inference strategies. By formalizing this interaction and introducing a practical algorithm that exploits it, our work offers a principled path toward more reliable and cost-effective alignment of black-box LLMs.
Figure 1: Inference-agnostic vs. inference-aware prompt optimization. The left side illustrates standard prompt optimization, which treats the inference strategy as fixed: a best prompt is selected during training and then used at inference with a predetermined number of samples, which can lead to misaligned outputs and high inference cost for some queries. The right side shows our inference-aware framework Iapo with the Psst algorithm, which conditions on user context such as budget and preferences, jointly selects the prompt and inference scale, and produces responses that better satisfy objectives and budget. Project page, code, and appendix are available online (https://iapo-aaai25.github.io/).
@@
In recent years, substantial efforts have been directed towards aligning large language models (LLMs) with human expectations in downstream tasks (mahmud2023explanation; Minaee et al. 2024). Many widely adopted alignment approaches—such as Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert 2025)—require access to model weights. This limitation has motivated a surge of interest in black-box alignment methods such as prompt optimization, which can align black-box models only through input manipulation (Ouyang et al. 2022; Zhou et al. 2022; Chang et al. 2024). Prompt optimization has demonstrated strong performance in both single-objective (Cheng et al. 2023; Trivedi et al. 2025) and multi-objective (Jafari et al. 2024; Zhao et al. 2025) settings. However, these methods remain agnostic to the inference strategy during deployment, potentially leading to suboptimal performance. In contrast, our approach explicitly captures the interdependence between inference-time strategies and prompt optimization.
@@
Recently, Shi et al. framed prompt optimization as a fixed-budget best-arm identification (BAI) problem. While effective under limited evaluation budgets, the method remains inference agnostic and was only explored in single-objective settings. Our work builds on this foundation in two key ways: (1) we introduce a contextual formulation that models user preferences over multiple objectives and associated computational costs; and (2) we incorporate inference-awareness to ensure alignment with the deployed inference strategy. To learn an optimal policy, we introduce a fixed-budget contextual BAI algorithm, Psst, inspired by Sequential Halving (SH) (Karnin et al. 2013). While SH was originally developed for the pure bandit setting, the Iapo framework features both inter-context full-information feedback and intra-context semi-bandit feedback. Psst leverages these structural properties to achieve more efficient optimization, extending beyond what standard SH can accommodate.
@@
Another relevant line of work focuses on inference-time alignment, where model outputs are improved during inference without modifying model parameters. Some of these methods, such as GenARM and DEAL (Xu et al. 2024; Huang et al. 2024), require access to model logits, limiting their applicability in black-box settings. In contrast, Best-of-N sampling (BoN) and Majority Voting (MV) methods operate purely on model outputs and have shown strong empirical gains by generating multiple candidates and selecting the best one (Krishna et al. 2022; OpenAI 2024; Yue et al. 2025). However, these approaches introduce a non-trivial computational cost, and to our knowledge, none of them explicitly optimize the trade-off between computational budget and output quality. Our initial experiments also indicate that inference scaling strategies have complex interactions with prompt design. Prompts that are optimized for single-shot decoding might not perform well with BoN or MV, and the reverse is also true. Therefore, an inference-aware prompt optimization framework is required.
@@
Finally, some white-box methods have recently integrated inference-awareness into the training process. Chow et al. (2025) proposed an inference-aware fine-tuning procedure that explicitly optimizes for exploration–exploitation trade-offs under BoN. Similarly, BOND (Sessa et al. 2024) and BonBon (Gui et al. 2024) aim to distill BoN policies into a single-pass decoding policy. While these approaches avoid the cost of sampling at inference time, they require full access to model parameters and do not generalize beyond BoN-style strategies. In contrast, our method complements inference-aware fine-tuning and is designed to operate in fully black-box settings.
Inference-Aware Prompt Optimization
@@
Let $\mathcal{X}$ be the set of user queries and $\mathcal{P}$ a finite prompt set. A pair $(x\in\mathcal{X},p\in\mathcal{P})$ is submitted to a frozen black-box LLM, which, under fixed decoding hyperparameters, generates $N\in[N_{\max}]\ (\text{i.e., }\{1,\dots,N_{\max}\})$ i.i.d. completions $\mathbf{y}_{1:N}=(y_{1},\dots,y_{N})$. $K$ bounded objectives (e.g. helpfulness, harmlessness, exact-match) score each completion $O_{k}:\mathcal{X}\times\mathcal{P}\times\mathcal{Y}\to[o_{k}^{\min},o_{k}^{\max}]$ where $\mathcal{Y}$ denotes the space of model completions. We also define the cost of producing a response as $\mathrm{Cost}(x,p,y_{i})$, a composite function that takes into account various computational factors such as token count, time, and energy. We add it as a (K+1)(K{+}1)-st objective $O_{K+1}=-\mathrm{Cost}(x,p,y_{i})$. An external entity supplies a context $c=(w_{1},\dots,w_{K+1})\in\mathcal{C}$, where every $w_{k}$ is chosen from a finite discrete domain. Given the above setup, we now formalize the inference strategies.
@@
$R_{x}^{\textsc{BoN}}(c,p,N)=\\ \underbrace{\max_{i\leq N}\sum_{k=1}^{K}w_{k}\,O_{k}(x,p,y_{i})}_{\text{task reward}}+\underbrace{w_{K+1}\sum_{i=1}^{N}O_{K+1}(x,p,y_{i})}_{\text{inference cost}}.$ (1)
@@
For query $x$, the pair $(p,N)$ yields i.i.d. completions $\mathbf{y}_{1:N}$ and extracted answers $\ell_{1:N}$. For each distinct answer $s$, define the vote count $n_{s}=\sum_{i=1}^{N}\mathds{1}[\ell_{i}=s]$, the maximum $n^{\star}=\max_{s}n_{s}$, and the tie multiplicity $t=\sum_{s}\mathds{1}[n_{s}=n^{\star}]$. MV predicts uniformly at random among the $t$ maximizers. With gold answer $\gamma$ and the success credit defined as $O_{1}(x,p,\mathbf{y}_{1:N})=\frac{\mathds{1}[\,n_{\gamma}=n^{\star}\,]}{t}$, we define MV utility as:
$R_{x}^{\textsc{MV}}(c,p,N)=\underbrace{w_{1}\,O_{1}(x,p,\mathbf{y}_{1:N})}_{\text{task reward}}+\underbrace{w_{2}\sum_{i=1}^{N}O_{2}(x,p,y_{i})}_{\text{inference cost}}.$ (2)
@@
(a) MATH, MV
(b) Helpful-Harmless, BoN (Prompt-A)
(c) Helpful-Harmless, BoN (Prompt-B)
Figure 2: Prompt–Inference Interdependence. (a) Accuracy under MV with Llama-3.3-70B-Instruct, showing prompt dominance shifts with budget (shaded). (b, c) Cost-adjusted reward under BoN decoding. Prompt and inference scales vary with user-specified trade-offs.
@@
Let an inference configuration be a tuple $g\in\mathcal{G}$ (e.g. temperature, top-$p$, max token). Then we define a set of arms $\mathcal{A}$ in Iapo as: $a\;=\;(p,g,N)\;\in\;\mathcal{A}:=\mathcal{P}\times\mathcal{G}\times[N_{\max}].$
@@
Given a dataset $\mathcal{X}$, context $c\in\mathcal{C}$, and aggregator $\alpha\in\{\textsc{BoN},\textsc{MV}\}$, the expected utility of arm $a$, i.e., the context-action value function or $Q$-function is defined as:
@@
Note that $R_{x}^{\alpha}(c,a)$ is a random variable. Now, let the context-optimal arm be $a^{\star}(c)=\arg\max_{a}Q^{\alpha}(c,a)$; hence the optimal Iapo policy is defined as: $\pi^{\star}(c)\;=\;a^{\star}(c),\forall\,c\in\mathcal{C}$.
In this paper, we adopt a train-then-deploy setup to learn the optimal Iapo policy. Given a total completion budget of $T$, at each round the learner may adaptively select a subset of arms. For any selected arm $a=(p,N)$, it samples a query $x\sim\mathcal{X}$, obtains $N$ completions, and observes raw reward vectors $\mathbf{o}_{i}\in\mathbb{R}^{K+1}$ for all $i\in[N]$. This repeats until the budget is exhausted, i.e., $\sum N=T$. After spending the entire budget, the learner returns a deployment policy $\pi_{T}$. The performance of this policy is evaluated by the Average Contextual Return:
$\text{ACR}(\pi_{T})\;=\;\mathbb{E}_{c\sim\mathcal{C}}\bigl[Q^{\alpha}(c,\pi_{T}(c))\bigr],$ (4)
@@
To see how the green and blue trends can emerge, consider the following example. Suppose that in a reasoning task evaluated with MV, Prompt 1 has a $40\%$ success rate on Query 1 and a $90\%$ success rate on Query 2, while Prompt 2 has a $62\%$ success rate on both queries. The single-shot success rate of a prompt is the average of its per-query success rates; under this metric, Prompt 1 is preferred over Prompt 2 ($0.65$ vs. $0.62$). Under MV with $N=10$, however, the success probability of a prompt on a query is the probability that a majority of its $N$ sampled responses are correct. In this setting, one can verify that the effective success rate of Prompt 1 drops to approximately $0.63$, whereas that of Prompt 2 increases to approximately $0.77$, so Prompt 2 becomes preferred. This example illustrates how increasing $N$ can change the relative ranking of prompts and produce the observed trends.
(b,c) Best-of-$\boldsymbol{N}$ on Helpful-Harmless. We evaluate two prompts (A and B, see appendix) on the Helpful-Harmless benchmark (Bai et al. 2022) using Best-of-N decoding for $N\leq 24$. Each curve corresponds to a different user-defined trade-off between helpfulness and harmlessness, plotting the cost-adjusted reward averaged over 1000 queries (see the appendix for details). The optimal choice of prompt (A vs. B) and sampling budget (N) is highly sensitive to these preferences. For example, the prompt A is strictly preferred when helpfulness is weighted more heavily.
Having established the need for inference-aware optimization, we now examine the precise conditions under which joint optimization becomes essential. We start by defining the Inference-Agnostic (IA) utility, which does not simulate inference scaling during training and instead optimizes the average utility achieved per prompt. More formally:
@@
$R_{x}^{\textsc{IA}}(c,a=(p,N))=\frac{1}{N}\textstyle\sum_{i=1}^{N}\sum_{k=1}^{K+1}w_{k}O_{k}(x,p,y_{i}).$ (5)
@@
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $\sigma R_{x}^{\textsc{IA}}(c,a),\sigma>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$\mathbb{E}_{x\sim\mathcal{X}}\bigl[\sigma R_{x}^{\textsc{IA}}(c,a)+\mu\bigr]=\sigma Q^{\textsc{IA}}(c,a)+\mu.\quad$
Figure 3: Expected utility ($w_{k+1}=0$) for MV (left) and BoN (right). MV shows a sharp performance drop when the correctness probability $\theta$ drops below 0.5, whereas BoN is strictly concave.
The above also highlights that affine aggregation significantly simplifies inference-aware optimization. For instance, in a regression task where the aggregated prediction is the mean of multiple numeric predictions and the reward is defined by the mean squared error (MSE), the resulting quantity can, under certain assumptions, become an affine transformation of the IA utility, eliminating the need to simulate inference scaling during training. However, common inference scaling strategies like BoN and MV generally do not admit such affine formulations. While they can sometimes be expressed as non-affine transformations of the IA—such as in the Bernoulli case with large $N$, where $R_{x}^{\textsc{IA}}(c,a)\approx\theta$—these are special cases (Figure 3). Hence, trying to determine the prompt based on $Q^{\textsc{IA}}$ for BoN or MV can result in misalignment. This motivates the next section, where we develop a training method that handles the general Iapo setting beyond the affine regime.
@@
In this section, we propose a fixed-budget arm elimination-based strategy for training policy $\pi_{T}$, called Psst (Prompt Scaling via Sequential Trimming). We then provide a theoretical analysis that establishes error guarantees for Psst under a finite inference budget. Finally, we introduce a practical approximation heuristic that reduces the training-time inference budget without significantly compromising performance in many practical settings.
@@
Classical arm-elimination methods such as Sequential Elimination (Even-Dar et al. 2006) and Sequential Halving (Karnin et al. 2013) follow a simple recipe: (i) split the elimination process into multiple rounds; (ii) in each round, allocate the round budget across the surviving arms; and (iii) trim a subset of arms at the end of the round based on their estimates. However, Iapo departs from pure BAI settings in the following ways:
@@
  • Cross‑context reuse. One pull of $(p,N)$ on query $x$ yields the completion set $\mathbf{y}_{1:N}$ and objective vector set $\mathbf{o}_{1:N}$ that can be used to estimate $R^{\alpha}_{x}(c,p,N)$ for all $c\in\mathcal{C}$.
  • Nested sample reuse across inference scales. Pulling a larger scale subsumes smaller ones: a pull of $(p,N_{i})$ produces $\big\lfloor N_{i}/N_{j}\big\rfloor$ i.i.d. samples for arm $(p,N_{j})$ by partitioning the $N_{i}$ draws into disjoint groups of size $N_{j}$ and then recomputing BoN/MV on each group.
A key consequence is that, for a prompt, the largest surviving scale drives the budget. Let $N^{(r)}_{\max}(p)=\max\{\,N:\ (p,N)\ \text{survives at the start of round }r\,\}.$ If we allocate $K$ pulls to $(p,N^{(r)}_{\max}(p))$ in round $r$, then every surviving arm $(p,N)$ with $N\leq N^{(r)}_{\max}(p)$ automatically receives at least $K$ effective samples by block reuse. Thus, an effective arm elimination strategy should exploit both (i) cross‑scale reuse and (ii) cross‑context reuse when estimating $Q$-function, while being aware of asymmetric cost.
@@
1:Context set $\mathcal{C}$, prompt set $\mathcal{P}$, $N_{\max}$, Scaling strategy $\alpha$, Query Dataset $\mathcal{X}_{\text{train}}$, total pull budget $T$;
@@
7:  $\mathcal{A}^{(r)}\leftarrow\{a:\exists c,\ \text{F}_{c,a}=\textbf{true}\}$
8:  $n_{r}\leftarrow\bigl\lfloor T/R\bigr\rfloor$
9:  $\lambda^{(r)}\leftarrow\textsc{Allocate}\bigl(\text{F},n_{r}\bigr)$
10:  $\mathcal{B}\leftarrow\{\}$
11:  for $a\in\mathcal{A}^{(r)}$ do
12:   for $i=1...\lambda^{(r)}(a)$ do
13:     Sample $x\sim\mathcal{X}_{\text{train}}$
14:     $\mathcal{B}\leftarrow\mathcal{B}\ \cup(x,a)$